5. You are the manager of a monopoly and your demand and cost functions are given by
P = 225 - 2Q
and
c(Q) = 1,100 + 3Q2
a. What price-quantity combination maximizes your firm’s profits?
1
Expert's answer
2019-11-18T09:49:15-0500

"P = 225 - 2Q"

"c(Q) = 1100 + 3Q^2"


"TR=P*Q"

"TR=Q(225-2Q)"

"TR=225Q-2Q^2"



"MR=\Delta TR/\Delta Q"

"MR=\Delta 225Q-2Q^2/\Delta Q"

"MR=225-4Q"


"MC=\Delta TC/\Delta Q"

"MC=\Delta 1100 + 3Q^2/\Delta Q"

"MC=6Q"


 Monopoly firm profit maximize is, 


"MR=MC"

"225-4Q=6Q"

"Q=22.5"



"P = 225 - 2Q"

"P = 225 - 2*22.5"

"P=180"


Firm profit maximize price "\boxed{180}" and quantity "\boxed{22.5}"